OpenStudy (anonymous):

Given f(x) = 4x + 1 all over 3, solve for f^-1(4). a. 2 b. 4 c. 6 d. 8

OpenStudy (austinl):

Okay, firstly. We need to find the inverse of f(x). Do you know how to do this?

OpenStudy (anonymous):

No, could you explain?

OpenStudy (austinl):

Okay, what you need to do first is replace the f(x) with a y. You should get an equation that looks kinda like this. \(y=4x+1\) Then you need to solve for x. When you have done this, you swap the places of y and x. Think you could do this for me?

OpenStudy (anonymous):

x = y -1 all over 4?

OpenStudy (austinl):

\(y=4x+1\) \(y-1=4x\) \(\displaystyle \frac{y-1}{4}=x\) Then you forgot the last step, you replace the x with the y. \(\displaystyle \frac{x-1}{4}=y\) Then we have, \(\displaystyle f(x)=\frac{x-1}{4}\) Do you know how to solve from there?

OpenStudy (austinl):

\(\displaystyle f^{-1}(x)=\frac{x-1}{4}\) My apologies, I used the wrong notation. My bad.

OpenStudy (anonymous):

I understand that part but how would I go from there?

OpenStudy (jdoe0001):

\(\bf f^{-1}({\color{red}{ x}})=\cfrac{{\color{red}{ x}}-1}{4}\qquad \qquad f^{-1}({\color{red}{ 4}})=\cfrac{{\color{red}{ 4}}-1}{4}\)

OpenStudy (anonymous):

Oh I see, so the solution would be 4?

OpenStudy (jdoe0001):

is it?

OpenStudy (austinl):

No, do like @jdoe0001 said \(\bf f^{-1}(\color{red}{4})\) This just means that you plug in 4 for all instances of x in the function.

OpenStudy (anonymous):

f^-1(4) = 3/4?

OpenStudy (austinl):

I would agree.

OpenStudy (anonymous):

I don't get where I am supposed to go from here.

OpenStudy (austinl):

Um... wait... those are your options above? They make no sense. None of them can be a possible answer. @jdoe0001 any thoughts?

OpenStudy (jdoe0001):

ohhh shoot.. I see the mistake.. tis was a typo I gather ok so \(\bf f(x)={\color{red}{ y}}=\cfrac{4{\color{blue}{ x}}+1}{3}\qquad inverse\implies {\color{blue}{ x}}=\cfrac{4{\color{red}{ y}}+1}{3}\\ \quad \\ \textit{solve for "y"}\)

OpenStudy (jdoe0001):

hmmm it says \(\bf f^{-1}(3)\ not \ f^{-1}(4)\) btw

OpenStudy (anonymous):

My sincere apologies!

OpenStudy (anonymous):

I just noticed

OpenStudy (austinl):

Son of a.... @FlowerOfLife you need to use proper formatting. IN THE BEGINNING \(\displaystyle \bf f(x)=\frac{4x+1}{3}\) That makes a difference.

OpenStudy (anonymous):

I'm so sorry, I'm not quite sure how to use proper formatting.

OpenStudy (jdoe0001):

@FlowerOfLife so solve for "y" multiply both sides by 3 then subtract 1 from both then divide by 4 both

OpenStudy (austinl):

If you aren't sure how to use \(\bf \LaTeX\), something as simple as f(x)=(4x+1)/3 is better than saying "all over" It is super easy to lose that when looking at a problem. And, not to nit-pick, you solve for x :P

OpenStudy (anonymous):

So then the answer still is not 4?

OpenStudy (austinl):

\(\displaystyle \bf f(x)=\frac{4x+1}{3}\) \(\displaystyle \bf y=\frac{4x+1}{3}\) \(\displaystyle \bf x=\frac{3y-1}{4}\) \(\displaystyle \bf f^{-1}(x)=\frac{3x-1}{4}\) \(\displaystyle \bf f^{-1}(\color{red}{3})=\frac{3(\color{red}{3})-1}{4}=\frac{9-1}{4}=\frac{8}{4}=2\)

OpenStudy (anonymous):

Again sorry for all the confusion u_u

OpenStudy (austinl):

No problem, glad we caught it and put you on the right path! Don't want you to learn the wrong stuff :P